My last question, please help
Write a quadrataic equation in the variable x having the given numbers as solutions. Type the equation in standard form ax^2+bx+c=0 where a,b, and c are integers(no fractions) solutions -2/3,7/3
Thank you

Answer by stanbon(75887)

Write a quadrataic equation in the variable x having the given numbers as solutions. Type the equation in standard form ax^2+bx+c=0 where a,b, and c are integers(no fractions) solutions -2/3,7/3
-----
If x = -2/3
then 3x+2 = 0
----
If x = 7/3
then 3x-7 = 0
----
Multiplying you get:
y = (3x+2)(3x-7)
y = 9x^2-15x-14
---------------------
Cheers,
Stan H.
=====================